OpenStudy (anonymous):

help solve please 29 and 7/15 - 16 and 5/9

OpenStudy (radar):

29 7/15 -16 5/9 ---------- 13 and some fraction. The fractions are 7/15-5/9

OpenStudy (radar):

Find the lowest common denominator for both 15 and 9. In this case I would start with 15 and go up in multiples until I found a value that contains 9 as a factor. 15 no, 30 no, 45 Bingo.

OpenStudy (radar):

45 is the LCD (Lowest Common Denominator. Convert the fractions to an equivalent fraction with a denominator of 45. 7/15 = 21/45 5/9 = 25/45 Did you follow that procedure with understanding?

OpenStudy (radar):

Back to the problem. We are to subtract 25/45 from 21/45 we are going to get a negative results. No problem, that is ok!!! 21/45-25/45=-4/45 Now are solution is 13 -4/45. To express this with an answer with a single sign we combine and get an answer of: 12 41/45

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks so much that helped alot

OpenStudy (radar):

Great, did you follow the 13 -4/45 becoming 12 41/45 ??

OpenStudy (radar):

13=12 45/45 -4/45 _______________ 12 41/45

OpenStudy (radar):

Just in case!

OpenStudy (anonymous):

yeah can you help me with this one its 4/2.4 = 6/x

OpenStudy (radar):

Is that decimal supposed to be in the denominator?

OpenStudy (anonymous):

yeah

OpenStudy (radar):

Ok Cross multiply getting 4x=(6)(2.4)

OpenStudy (radar):

x=(6)(2.4)/6

OpenStudy (radar):

You can solve. I gotta run. Birthday part for great grand daughter.

OpenStudy (anonymous):

ok thanks have fun

OpenStudy (radar):

I made a typo should be x=(6)(2.4)/4